OpenStudy (anonymous):

if F(x)=tanx, then fprime(pi/6)=

OpenStudy (anonymous):

Have you differentiated F(x) and then sub'ed \(\frac{\pi}{6}\) into it?

OpenStudy (anonymous):

Do you know the derivative of tan(x) ??

OpenStudy (anonymous):

i kno the differvative is secxtanx but i dont know where to go frm there :/

OpenStudy (anonymous):

Yes, than just replace x by pi/6..

OpenStudy (anonymous):

(2\[\sqrt{3}\]/3 and \[\sqrt{3}\]/3

OpenStudy (anonymous):

\[F'(\frac{\pi}{6}) = \sec(\frac{\pi}{6}) \cdot \tan(\frac{\pi}{6})\]

OpenStudy (anonymous):

igot tht part

OpenStudy (anonymous):

What is the value of : \[\sec (\frac{\pi}{6})\]

OpenStudy (anonymous):

2root3/3

OpenStudy (anonymous):

Yes and what is for tan(pi/6) = ??

OpenStudy (anonymous):

root3/3

OpenStudy (anonymous):

Wait, derivative of tan(x) isn't it sec^2(x) ??

OpenStudy (anonymous):

its not ?

OpenStudy (anonymous):

\[\frac{d}{dx} \tan(x) = \sec^2(x)\]

OpenStudy (anonymous):

thts y im getting the wrong answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You were doing the derivative of sec(x)..

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\frac{d}{dx}\sec(x) = \sec(x) \cdot \tan(x)\] It is okay and sorry from my side too, I also got confused.. And you are welcome dear..

OpenStudy (anonymous):

ok idk how to combine the 2 radicals

OpenStudy (anonymous):

Where ??

OpenStudy (anonymous):

sec^2

OpenStudy (anonymous):

\[F'(\frac{\pi}{6}) = \sec^2(\frac{\pi}{6})\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\implies \frac{2 \sqrt{3}}{3} \times \frac{2\sqrt{3}}{3} \implies ??\]

OpenStudy (anonymous):

This is simple multiplication and division now...

OpenStudy (anonymous):

i know but im not getting the right answer

OpenStudy (anonymous):

4

OpenStudy (anonymous):

What are you getting tell me first...

OpenStudy (anonymous):

Its not 4... You are missing something.

OpenStudy (anonymous):

\[\sqrt{3} \times \sqrt{3} = ??\]

OpenStudy (anonymous):

2 times 2=4 root3 times root 3= 3 4 times 3 =12 12 divide3 =4

OpenStudy (anonymous):

12 divided by 3 ??? Are you sure in denominator it is 3 ??

OpenStudy (anonymous):

im not too good with my unit circle

OpenStudy (anonymous):

Dear, what will you get when you will square 3 ??

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Check the colored things here : \[\frac{2 \sqrt{3}}{\color{red}{3}} \times \frac{2 \sqrt{3}}{\color{red}{3}} = ??\]

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Yes it is 12 divided by 9 and not 3.. So what will you get now ??

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

i was thinking addition

OpenStudy (anonymous):

my mistake...srry if i took up yhur time

OpenStudy (anonymous):

It is okay but be careful from next time... These are called "Silly Mistakes..." Ok, now concentrate more on new problems...

OpenStudy (anonymous):

It is okay..

OpenStudy (anonymous):

thanks a lot :)))